OpenStudy (anonymous):

Geometry help please! The longest side of a triangle is 3 greater than the longest side of a similar triangle. If the ratio of similitude is 5:4, find the lengths of the longest side of each triangle.

OpenStudy (anonymous):

\[\frac{ x+3 }{ x } = \frac{ 5 }{ 4 }\]

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

x = 12

OpenStudy (anonymous):

^

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So the longest sides are 12 and 15.

OpenStudy (anonymous):

awesome. Thank you so much.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

haha i have one more coming and thats it